a) Determine the point estimate of the population mean and margin of error for the confidence interval Lower bound is 20, upper









b)In a poll, 51% of the people polled answered yes to the question Are you in favor of the death penalty for a person convicte
Answer #1

Solution :

\bar x  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = \bar x = (20+24) /2 =22

point estimate=Sample mean = \bar x   =22

Margin of error = E = Upper confidence interval - \bar x   = 24-22=2

Margin of error = E =2

(B)

Solution :

Given that,

\hat p= 0.51

1 - \hat p = 1 - 0.51 = 0.49

margin of error = E = 2% = 0.02

At 94% confidence level the z is ,

Z\alpha/2 = Z0.03 = 1.88 ( Using z table )

Sample size = n = (Z\alpha/2 / E)2 * \hat p * (1 - \hat p )

= (1.88 / 0.02)2 * 0.51 * 0.49

= 2208.1164

Sample size =2209